# Gravitational Wave emission in Binary Neutron Star early post-merger within a dark environment

### v1

@@ III Spectral properties
where $t_{w}$ is the early-postmeger window time $\Delta t$ that we defined for each signal, see Fig. (2). The transform is taken over this polarization mode due to the approximation to the principal mode $h_{+}\simeq h^{22}$.
@@
In Fig. (3) we plot our results for the three data sets THC005, THC0032, and THC0040 in the left, center, and right columns, respectively. We depict with a thin red line the full post-merger stage of the data and with thick orange, green, and blue lines the early post-merger interval $\Delta t$ associated with each dataset utilized. We set $\beta_{2}=0$ in this panel for all studied cases. The orange line shows the DM free case, while the orange and blue set $f_{\chi}=1\%$ and $f_{\chi}=5\%$ DM mass fraction, respectively. For the upper panel $\beta_{1}=0.5\times 10^{16}$ $\rm(cm/s)^{2}$, middle $\beta_{1}=0.25\times 10^{16}$ $\rm(cm/s)^{2}$ and lower $\beta_{1}=0$. In order to see how feasible a future detection may be we superimpose the sensitivity curves of the Advanced LIGO, ET, NEMO and Cosmic Explorer detectors. We obtain $f_{\rm max}$=2.3426, 2.9608 and 2.6667 kHz for THC0005, THC0032,THC0040 cases, respectively. Note that the DM influence related to mass fraction is expected to be small in $\Lambda$ as already shown in Fig. (5) in our previous work [4]. For our set of EoS the nuclear matter incompressibility $K\in[242.7,245.5]$ MeV at saturation density, thus to this respect we are showing cases with a small stiffness spread. We nevertheless see that as shown in [4] small compactness stars are more sensitive to a $f_{\chi}\neq 0$ fraction and thus more likely to show deviations of the DM free $f_{\rm max}$ value.
THC0005 $\beta_{2}$ $[\rm 10^{8}\,cm/s]$ $\mu_{\chi}$ $[10^{30}\rm g/cms]$
$f_{\chi}=0.01$ $\rho_{\chi,14}=0.0463$ $10^{2}$ $0.106$
$10^{3}$ $1.060$
$10^{4}$ $10.60$
$f_{\chi}=0.05$ $\rho_{\chi,14}=0.232$ $10^{2}$ $0.528$
$10^{3}$ $5.28$
$10^{4}$ $52.8$
THC0032
$f_{\chi}=0.01$ $\rho_{\chi,14}=0.0462$ $10^{2}$ $0.158$
$10^{3}$ $1.580$
$10^{4}$ $15.80$
$f_{\chi}=0.05$ $\rho_{\chi,14}=0.231$ $10^{2}$ $0.789$
$10^{3}$ $7.890$
$10^{4}$ $78.90$
THC0040
$f_{\chi}=0.01$ $\rho_{\chi,14}=0.0484$ $10^{2}$ $0.126$
$10^{3}$ $1.260$
$10^{4}$ $12.60$
$f_{\chi}=0.05$ $\rho_{\chi,14}=0.242$ $10^{2}$ $0.528$
$10^{3}$ $5.28$
$10^{4}$ $52.8$
Table 3: The table shows the values of DM parameters $f_{\chi}$ and $\beta_{2}$ used in Fig. (4), along with their reference values for density $\rho_{\chi}$ in units of $10^{14}$ $\rm g/cm^{3}$ and viscosity $\mu_{\chi}$ in the cgs unit system.
Upon analyzing the frequency spectrum families obtained from our model, it becomes evident how accurately the model replicates the principal peak $f_{\rm max}$, which characterizes the post-merger signal, across the entire temporal interval of the simulated signal. This feature is expected, as this peak is commonly attributed to the inherent quadrupolar oscillation of the system. Furthermore, the model successfully captures a secondary, lower frequency peak, $f_{1}$, at low frequencies associated with a non-linear coupling between the quadrupolar oscillation modes and the quasi-radial modes. It should be noted, in addition, that the primary contributions belong to kHz frequencies, where the peaks appear to originate almost entirely during the early post-merger stage, as illustrated in [1] for the peak at lower frequencies. Similarly, concerning the largest amplitude peak, it has been observed to grow rapidly in the early post-merger stage and then remain almost unchanged throughout the remainder of the post-merger period [57].
@@
Figure 4: Panel represents the spectral density $\sqrt{S_{h}(f)}=2\tilde{h}(f)f^{1/2}$ computed for the full post-merger data (thin red line) and for the early post-merger interval after the fitting procedure (thick colored lines) obtained from variation of DM parameters ($f_{\chi}$, $\beta_{2}$) with $\beta_{1}=0$. We also plot sensitivity curves of the NEMO, CE, ET and aLIGO detectors (black and grey curves). Data sets THC005, THC0032, and THC0040 appear in the left, center, right columns, respectively. Distance is fixed to $d=50$ Mpc.
@@
In the following, we discuss the obtained results for the PSD. We first consider that the environmental DM causes dissipation in the system that is constant and proportional to linear DM density. This behavior is sometimes referred to in the literature as Coulomb damping [58]. In this case, there is only a non-zero $\beta_{n}$ i.e. $\beta_{1}\neq 0$ in Section II and we fix $\beta_{2}=0$. For decreasing values of $\beta_{1}$ characterizing DM behavior we show (left, center and right panels) the PSD for the THC005, THC0032, and THC0040 cases (thick colored lines) considered in Fig. (3). We set $d=50$ Mpc. The amplitude of the frequency peaks exhibits more pronounced deviations from the DM-free case ($f_{\chi}=\beta_{n}=0$). This is especially relevant in the right end of the $f_{\chi}\in[0.01,0.05]$ interval. Importantly, the decrease in the amplitudes in $\sqrt{S_{h}(f)}=2\tilde{h}(f)f^{\frac{1}{2}}$ from DM dissipation may result in the post-merger region of the event no longer being resolved by the detector at that distance. Consequently, events could be detected at distances where the inspiral phase is well-determined, but not that of the post-merger phase. In other words, detections of BNS mergers in environments containing modest quantities of viscous DM might exhibit GW emission where the typical post-merger early-phase behavior is not discernible, resembling, instead, a rapid decay toward a BH, for example. As shown in the lower panels, when the fraction of DM remains below the fiducial $\sim 5\%$ threshold and it is assumed a non-viscous behavior, the spectral density curve exhibits only marginal deviations from the DM free pattern. This observation suggests that the introduction of a limited amount of non-viscous matter does not significantly disrupt the expected spectral characteristics.
In Fig.(4) we analyze the so-called viscous damping and set $\beta_{1}\equiv 0$ to examine the frequency spectrum for different values of $\beta_{2}\neq 0$, using the same mass fractions $f_{\chi}=0,0.01,0.05$ as in the previous Fig. (3). Following the same logic as in section (II.1.4) and setting the values of $\beta_{2}=0,10^{2},10^{3},10^{4}$ (in $\rm 10^{8}cm/s$ units) the resulting viscosity values are presented in the Table (3). Note that the range in $\beta_{2}$ values is larger as it corresponds to tiny $\sim v_{r}^{2}$ dependences.
Let us now consider in more detail the two scenarios for the environmental conditions typical from merging NS [38]. For the cases analyzed we will set DM density somewhat lower than baryonic matter density. In units of nuclear saturation density $\rho_{\rm sat}\sim 2\times 10^{14}$ $\rm g/cm^{3}$, DM density, mass fraction values are displayed in Table (3) along with viscous parameters and its effective conversion to more familiar $\rm g/cms$ cgs units. Thus for $\rho_{\chi,14}=0.1,0.01$ in units of $10^{14}$ $\rm g/cm^{3}$ the conversion dictates also extreme values for $\mu_{\chi}\sim[0.1,100]\times 10^{30}$ $\rm g/cms$.
@@
In this case of viscous damping, the characteristic peaks of the PSD curve disappear. In Fig. (4), it can be observed that for the case with $f_{\chi}=0.05$ and $\beta_{2}=10^{4}$ (in $\rm 10^{8}\,cm/s$ units), the signal vanished entirely, in contrast with the behavior obtained for DM with Coulomb-type damping, where the signal peak amplitude decreases while maintaining its characteristic shape.
@@
In order to discuss the prospects of detectability we proceed to quantify the damping or attenuation in the maximum peak amplitude $f_{\text{max}}$ in the PSD due to the existence of a DM polluted environment. For this we calculate the effective signal-to-noise ratio (eSNR) as done in [1]. In our case this could be defined as
@@
The eSNR, as a function of the DM phase space $(\beta_{1},f_{\chi})$, constitutes a useful and straightforward tool that allows quantifying the influence of DM on the attenuation of peak amplitudes and the rest of the modes. Note that, in principle, this quantity is generally frequency dependent, eSNR(f). Considering the lower bound given by the sensitivity curve of aLIGO (indicated by the black dashed lines in Fig. (3), we observe in Fig. (5) the eSNR isocurves dependence on ($\beta_{1},f_{\chi}$) for the THC0032 case (bottom) and in more detail the PSD (top) for selected $\beta_{1}=0.33,0.37,0.4,0.42,0.5$ in $\rm 10^{16}(cm/s)^{2}$ units and $f_{\chi}=0.045$ values. Since $\beta_{2}$ is so efficient in the damping behaviour we choose for the sake of illustration, instead, the $\beta_{1}$ Coulomb damping. For all cases, we assumed GW detections with optimal orientation at 50 Mpc. At this point we remark that the spectral analysis we perform is conducted within a $\Delta t$ temporal window and consequently, the depicted PSD reflects only the frequencies that influence the signal’s power within this interval.
@@
We rely on previous results, which suggest that the main peak is predominantly associated with the early post-merger phase. During this interval, kHz frequencies are more prominent. However, when DM is introduced, the main peak amplitude decreases, reducing the contribution of these frequencies to the signal. From inspection of the eSNR in the bottom panel we see that there is an enhancement, breaking the attenuation tendency, of the peak amplitude for a band (upper right corner). This is due to the fact that in this band the peak amplitude shifts towards lower frequency $\sim 500$ Hz (thus reversing the eSNR ratio decrease) and the original amplitude around $\sim 3$ kHz falls within aLIGO sensitivity. The detail of this behaviour is illustrated in the top panel where for $\beta_{1}\in[0.33,0.5]$ and $f_{\chi}=0.045$ the monotony of the main amplitude ratio changes.
@@ IV Conclusions
We have developed a semi-analytical mechanical Lagrangian model inspired by [1, 17] where an effective elastic potential allows to simulate the dynamics in binary NS mergers in the presence of a DM-polluted environment. This model successfully reproduces the emitted GW signal in a time window for selected early post-merger instances of the CoRe simulation database with no DM. We also analyzed the resulting frequency spectrum and confirmed that the characteristic large amplitude mode, $f_{\rm max}$ appearing in the so-called quasi-universal relations, can be recovered in this scenario. By making relatively loose assumptions about the dimensions of the debris disk formed during the early post-merger phase and leveraging the numerical solution results of the model, we obtained values close to those predicted by numerical relativity simulations regarding baryonic density and viscosity.
When the initial (DM free) model is supplemented in a minimal procedure with a non-zero viscous mass fraction of DM in the BNS environment the qualitative behaviour of the GW changes exhibiting a pronounced damping. By analyzing the shape of the frequency spectrum, particularly its characteristic large amplitude peak, we tested the implications of adding up to realistic $5\%$ of matter from the environment. Three fundamental cases were analyzed: matter that does not dissipate the system’s energy, and matter that causes dissipation, where we considered two types of dissipation—Coulomb and viscous. For the inviscid matter, we found negligible changes in the PSD up to $5\%$ DM analyzed. For Coulomb/viscous damping, we found that the amplitude of the main peak decreases appreciably, not resulting shifted from the $\sim 3$ kHz value except for some DM parameter region. However frequency dependent DM-induced attenuation does not affect low frequency models as much. This is suggesting that matter with this type of behavior could lead to detections where this peak is not resolved by interferometers and could mimic instead a prompt collapse to a more compact object. When considering the effective SNR we find aLIGO yields a lower bound and could detect the possible effects of damping in maximum amplitude mode yielding a non-monotonic tendency for a subset of DM parameters . There is thus a possible interplay among (ordinary) matter effects and those of viscous DM that should not be overlooked. Further work is necessary to investigate these aspects and elucidate the importance of environmental effects and its imprint in waveforms.
@@ Acknowledgements
We acknowledge useful discussion from D. Radice, A. Perego, I. de Martino, A. de la Cruz-Dombriz. R. Della Monica and P. Char. We also acknowledge support from Spanish Ministry of Science PID2022-137887NB-10 and RED2022-134411-T projects and European Union’s HORIZON MSCA-2022-PF-01-01, project ProMatEx-NS. D. Suarez-Fontanella gratefully acknowledges the financial support provided by the Banco Santander International Doctoral Scholarship. D. Barba-González acknowledges support from a Ph.D. Fellowship funded by Consejería de Educación de la Junta de Castilla y León and European Social Fund.We acknowledge use of Spanish RES resources.

### v2

@@ III Dark Matter imprints on the spectral distribution
where $t_{w}$ is the early-postmerger window time $\Delta t$ that we defined for each signal, see Fig. (2). Note that from construction this will not match the full Fourier transform of the whole simulation data even in absence of DM but we take this as a benchmark to compare our findings. The transform is taken over this polarization mode due to the approximation to the principal mode $h_{+}\simeq h^{22}$.
@@
Figure 4: Panel represents the spectral density $\sqrt{S_{h}(f)}=2\tilde{h}(f)f^{1/2}$ computed for the full post-merger data (thin red line) and for the early post-merger interval after the fitting procedure (thick colored lines) obtained from variation of DM parameters ($f_{\chi}$, $\beta_{2}$) with $\beta_{1}=0$. We also plot sensitivity curves of the NEMO, CE, ET and aLIGO detectors (black and grey curves). Data sets THC0005, THC0032, and THC0040 appear in the left, center, right columns, respectively. Distance is fixed to $d=50$ Mpc.
In Fig. (3) we plot our results for the three data sets THC0005, THC0032, and THC0040 in the left, center, and right columns, respectively. We depict with a thin red line the full post-merger stage of the data and with thick orange, green, and blue lines the early post-merger interval $\Delta t$ associated with each dataset utilized. We set $\beta_{2}=0$ in this panel for all studied cases. The orange line shows the DM free case, while the orange and blue set $f_{\chi}=1\%$ and $f_{\chi}=5\%$ DM mass fraction, respectively. For the upper panel $\beta_{1}=0.5\times 10^{16}$ $\rm(cm/s)^{2}$, middle $\beta_{1}=0.25\times 10^{16}$ $\rm(cm/s)^{2}$ and lower $\beta_{1}=0$. In order to see how feasible a future detection may be we superimpose the sensitivity curves of the Advanced LIGO, ET, NEMO and Cosmic Explorer detectors. We obtain $f_{\rm max}$=2.3426, 2.9608 and 2.6667 kHz for THC0005, THC0032,THC0040 cases, respectively. For our set of EoS the nuclear matter incompressibility $K\in[242.7,245.5]$ MeV at saturation density, thus to this respect we are showing cases with a small stiffness spread. We nevertheless see by simple inspection that small compactness stars are more sensitive to a $f_{\chi}\neq 0$ fraction and thus more likely to show deviations of the DM free $f_{\rm max}$ value [4].
Upon analyzing the frequency spectrum obtained from our model, it becomes evident that for the nearly asymmetric cases we consider this model reasonably provides the principal peak $f_{\rm max}$, which characterizes the post-merger signal, across the entire temporal interval of the simulated signal.
Further comparison shows the model captures a secondary, lower frequency peak at low frequencies associated with a non-linear coupling between the quadrupolar oscillation modes and the quasi-radial modes. It should be noted, in addition, that the primary contributions belong to kHz frequencies, where the peaks appear to originate almost entirely during the early post-merger stage, as illustrated in [1] for the peak at lower frequencies. Similarly, concerning the largest amplitude peak, it has been observed to grow rapidly in the early post-merger stage and then remain almost unchanged throughout the remainder of the post-merger period [57].
@@
In the following, we discuss the obtained results for the PSD. We first consider that the environmental DM causes dissipation in the system that is constant and proportional to linear DM density. This behavior is sometimes referred to in the literature as Coulomb damping [58]. In this case, there is only a non-zero $\beta_{n}$ i.e. $\beta_{1}\neq 0$ in Section II and we fix $\beta_{2}=0$. For decreasing values of $\beta_{1}$ characterizing DM behavior we show (left, center and right panels) the PSD for the THC0005, THC0032, and THC0040 cases (thick colored lines) considered in Fig. (3). We set $d=50$ Mpc. The amplitude of the frequency peaks exhibits more pronounced deviations from the DM-free case ($f_{\chi}=\beta_{n}=0$). This is especially relevant in the right end of the $f_{\chi}\in[0.01,0.05]$ interval. Importantly, the decrease in the amplitudes in $\sqrt{S_{h}(f)}=2\tilde{h}(f)f^{\frac{1}{2}}$ from DM dissipation may result in the post-merger region of the event no longer being resolved by the detector at that distance. Consequently, events could be detected at distances where the inspiral phase is well-determined, but not that of the post-merger phase. In other words, detections of BNS mergers in environments containing modest quantities of viscous DM might exhibit GW emission where the typical post-merger early-phase behavior is not discernible, resembling, instead, a rapid decay toward a BH, for example. As shown in the lower panels, when the fraction of DM remains below the fiducial $\sim 5\%$ threshold and it is assumed a non-viscous behavior, the spectral density curve exhibits only marginal deviations from the DM free pattern. This observation suggests that the introduction of a limited amount of non-viscous matter does not significantly disrupt the expected spectral characteristics.
@@
In Fig.(4) we analyze cases with $\beta_{1}\equiv 0$ to examine the frequency spectrum for different values of $\beta_{2}\neq 0$, using the same mass fractions $f_{\chi}=0,0.01,0.05$ as in the previous Fig. (3). We can observe that the characteristic peaks of the PSD curve disappear. For the case with $f_{\chi}=0.05$ and $\beta_{2}=10^{4}$ (in $\rm 10^{8}\,cm/s$ units), the signal vanished entirely, in contrast with the behavior obtained for DM with Coulomb-type damping, where the signal peak amplitude decreases while maintaining its characteristic shape.
@@
In order to discuss the prospects of detectability we proceed to examine the damping or attenuation in the maximum peak amplitude $f_{\text{max}}$ in the PSD due to the existence of a DM polluted environment. For this we calculate the effective signal-to-noise ratio (eSNR) as done in [1]. In our case this could be defined as
@@
The eSNR, as a function of the DM phase space $(\beta_{1},f_{\chi})$, constitutes a useful and straightforward tool that allows sizing the influence of DM on the attenuation of peak amplitudes and the rest of the modes. Note that, in principle, this quantity is frequency dependent, eSNR(f). Considering the lower bound given by the sensitivity curve of aLIGO (indicated by the black dashed lines in Fig. (3), we observe in Fig. (5) the eSNR isocurves dependence on ($\beta_{1},f_{\chi}$) for the THC0032 case (bottom) and in more detail the PSD (top) for selected $\beta_{1}=0.33,0.37,0.4,0.42,0.5$ in $\rm 10^{16}(cm/s)^{2}$ units and $f_{\chi}=0.045$ values. Since $\beta_{2}$ largely impacts the damping behavior we choose for the sake of illustration, instead, the $\beta_{1}$ damping. For all cases, we assumed GW detections with optimal orientation at 50 Mpc. At this point we remark that the spectral analysis we perform is conducted within a $\Delta t$ temporal window and consequently, the depicted PSD reflects only the frequencies that influence the signal’s power within this interval.
@@
We rely on previous results, which suggest that the main peak is predominantly associated with the early post-merger phase. During this interval, kHz frequencies are more prominent. However, when DM is introduced, the main peak amplitude decreases, reducing the contribution of these frequencies to the signal. From inspection of the eSNR in the bottom panel we see that there is an enhancement, breaking the attenuation tendency, of the peak amplitude for a band (upper right corner). This is due to the fact that in this band the peak amplitude shifts towards lower frequency $\sim 500$ Hz (thus reversing the eSNR ratio decrease) and the original amplitude around $\sim 3$ kHz falls within aLIGO sensitivity. The detail of this behavior is illustrated in the top panel where for $\beta_{1}\in[0.33,0.5]$ and $f_{\chi}=0.045$ the monotony of the main amplitude ratio changes.
@@ IV Conclusions
Taking an effective Lagrangian model partly inspired by [1, 16] we qualitatively explore the dynamics of nearly symmetric binary NS mergers in the presence of a DM-polluted environment. As a benchmark, this model allows a reasonable reproduction of the emitted GW signal in a time window for selected early post-merger instances of the CoRe simulation database with ordinary matter. We use MCMC techniques fitting a n-tuple parameter for each selected BNS configuration. We also obtained the resulting frequency spectrum and confirmed that the characteristic large amplitude mode, $f_{\rm max}$ can be recovered in this scenario where $q\sim 1$. When the initial (DM free) model is supplemented in a minimal procedure with a non-zero viscous mass fraction of DM in the BNS environment, the qualitative behavior of the GW changes exhibiting a pronounced damping. By analyzing the shape of the frequency spectrum, particularly its characteristic large amplitude peak, we tested the implications of adding up to realistic $5\%$ of matter from the environment. We analyzed cases where DM matter was present but not inducing dissipative effects or it produced dissipation instead. For the inviscid matter, we found negligible changes in the PSD up to $5\%$ DM analyzed. For Coulomb/viscous damping, we found that the amplitude of the main peak decreased appreciably, not resulting shifted from the $\sim 3$ kHz value except for a restricted DM parameter region. However frequency dependent DM-induced attenuation does not affect low frequency models as much. This is suggesting that matter with this type of behavior may lead to detections where this peak is not resolved by interferometers and could mimic instead a prompt collapse to a more compact object. When considering the effective SNR we find aLIGO yields a lower bound and could detect the possible effects of damping in maximum amplitude mode yielding a non-monotonic tendency for a subset of DM parameters. There is thus a possible interplay among (ordinary) matter effects and those of viscous DM that should not be overlooked as pointed out by recent works [39]. Further work is necessary to investigate these aspects and elucidate the importance of environmental effects and its imprint in waveforms.
@@ Acknowledgements
We acknowledge useful discussion from D. Radice, A. Perego, I. de Martino, A. de la Cruz-Dombriz. R. Della Monica and P. Char. We also acknowledge support from Spanish Ministry of Science PID2022-137887NB-10 and RED2022-134411-T projects, Junta de Castilla y León projects SA101P24, SA091P24 and European Union’s HORIZON MSCA-2022-PF-01-01, project ProMatEx-NS. This publication is based upon work from the COST Action “COSMIC WISPers” (CA21106) supported by COST (European Cooperation in Science and Technology). D. Suarez-Fontanella gratefully acknowledges the financial support provided by the Banco Santander International Doctoral Scholarship. D. Barba-González acknowledges support from a Ph.D. Fellowship funded by Consejería de Educación de la Junta de Castilla y León and European Social Fund.We acknowledge use of Spanish RES resources.